I have no idea how to go about these questions. Could someone please help.

1. Use your estimated molecular mass to calculate the number of amino acids in the human serum protein.

2. Using your estimated molecular mass, calculate the molarity of the albumin in the human serum sample.

MW (kDa)

Log MW

Distance (units unknown)

250

2.397940009

18

150

2.176091259

76

100

2

137

75

1.875061263

187

50

1.698970004

265

37

1.568201724

327

25

1.397940009

418

20

1.301029996

455

15

1.176091259

537

10

1

600

Human Blood serum Well Lane

Distance (unknown units)

Estimated Molecular Mass of HBS (kDa)

Albumin Lane 1*

289

48.9

Albumin Lane 2

288

49.1

Albumin Lane 4

274

52.9
Answer #1

1. Average molecular weight of one amino acid = 110Da or 0.11kDa

Molecular weight (MW) of albumin = 48.9 kDa

Number of amino acids in albumin = 48.9/0.11 = 445 amino acids (approximately)

2. Molarity of protein in a sample = Concentration of protein per mL/ MW of protein in kDa

Concentration of albumin in a healthy individual = 5 g/dL or 0.05 g/ml or 50mg/ml

MW of albumin = 48.9kDa

Molarity = 50/48.9 = 1.02mM